11. What is the hourly flow rate for 250 mL of normal saline to be administered over 2 hours? _____ mL/h

12. A patient is to receive 1000 mL of D5W intravenously over 6 hours. The tubing set has a drop factor of 15 gtt/mL. How many drops per minute are needed to infuse this fluid in the prescribed time? _____ gtt per minute

13. What is 12% of 60?

a. 5

b. 7.2

c. 10.6

d. 12.8

14. The nurse is caring for a client on the medical surgical floor. The client consumed 150 mL of apple juice, 1/2 pint of milk, 1/2 cup of coffee, 3 ounces of water and 1 bowl of oatmeal without milk. Calculate the client’s intake in mL. ________mL.

15. A patient who weighs 168 lb is prescribed 15 mg of theophylline per kilogram of body weight intravenously as a loading dose. The drug is available at 250 mg/10 mL. How many milliliters should you prepare for this patient? Round to the nearest 10th of a mL. ___________mL

16. One quart equals _____cups

a. 3

b. 1

c. 5

d. 4

e. 6

f. 2

17. 2600gm= ? kg

a. 260 kg

b. 0.026 kg

c. 0.26 kg

d. 2.6kg

18. Safe drug administration requires that the individual giving a drug be knowledgeable about which drug features except?

a. Abnormal reactions

b. Actions

c. Follow-up care

d. Side effects

e. Purpose

f. Brand name 1

9. What technique should you use when measuring a liquid drug in a medicine cup to ensure an accurate dose?

a. Use a syringe to draw up the drug and then place it in the medicine cup.

b. Hold the cup above eye level while pouring the drug.

c. Hold the cup below eye level while pouring the drug.

d. Place the cup on a table and bend to look straight at the mark
Answer #1

11. The flow rate for 2 hours is 250ml of normal saline

Flow rate per hour = Total volume \div time in hours

                              = 250 \div 2 = 125 ml per hour

12. Drops per minute = Total volume \times drop factor \div time in minutes

                                   = 1000 ml \times 15 gtt \div 360 minutes

                                   = 41.66 gtt per min

13. 12% of 60 = 12 \div 100 \times 60

                       = 7.2

Option b is the correct answer.

14. The client consumed 150 mL of apple juice

1/2 pint of milk

1/2 cup of coffee

3 ounces of water

1 bowl of oatmeal without milk

The client’s intake is 150 + 236.5 + 118 + 88.7 = 593.2 ml or 593 ml